# AutoScale: Automatic Prediction of Compute-optimal Data Composition for Training LLMs

### v2

@@ Abstract
Domain reweighting is an emerging research area aimed at adjusting the relative weights of different data sources to improve the effectiveness and efficiency of language model pre-training. This paper demonstrates that the optimal composition of training data from different domains is scale-dependent, challenging the existing practice of determining optimal mixtures through small-scale experiments and directly applying them at larger scales. We derive an analytical model for the dependence of optimal weights on data scale and introduce AutoScale, a novel, practical approach for optimizing data compositions at potentially large training data scales. AutoScale first uses a principled optimization framework to find optimal compositions at smaller, feasible scales, then predicts optimal compositions at larger scales using our derived model. Our evaluation on GPT-2 Large and BERT pre-training demonstrates AutoScale’s effectiveness in improving training convergence and downstream performance. Particularly, for GPT-2 Large on RedPajama, AutoScale decreases validation perplexity 28% faster than baselines, with up to 38% speed-up over unweighted training, achieving the best performance across downstream tasks. This work provides insights into the varying benefits of data sources across training scales for language models, contributing to the burgeoning research on scale-dependent data curation. Code is open-sourced [^1] https://github.com/feiyang-k/AutoScale.
@@ 1 Introduction
Large language models (LLMs) are pre-trained on vast datasets sourced from diverse domains. However, the immense computational demands of this process, coupled with limited resources, create a pressing need to enhance the effectiveness and efficiency of pre-training. A promising approach to address this challenge is through domain reweighting—adjusting the relative proportions of data from different sources (Xie et al., 2024; Chen et al., 2024; Albalak et al., 2023; Soldaini et al., 2024; Fan et al., 2023; Ye et al., 2024).
Showing encouraging potentials, though, current implementation techniques face significant limitations. A prevailing technique is to first optimize data composition for a smaller proxy model and at a smaller data scale (Xie et al., 2024; Fan et al., 2023; Ye et al., 2024; Liu et al., 2024). Yet, this optimization often employs alternative objectives that may not align with the primary goal of minimizing evaluation loss. Moreover, the optimized weights are directly applied to training the target model on much larger data scales, implicitly assuming that the "optimal data composition" remains constant across data scales. This assumption of scale-invariance, however, may not hold in practice, potentially leading to suboptimal performance when scaling up. While research has scale-dependent data selection at the individual data point level for vision models (Sorscher et al., 2022; Goyal et al., ), it remains unclear whether this scale dependence applies to domain-level optimization, or how such scaling behavior might manifest in language models.
In parallel, an increasingly popular practice (Mehta et al., 2024) is to directly adopt domain weights that were designed for training previous models, such as those used for LLaMA (Touvron et al., 2023). However, these weights are optimized for specific applications that may differ from the desired use case of the target model. Given the limitations of these approaches, many in the industry still rely on heuristics for mixing domain data (McKinzie et al., 2024; Mehta et al., 2024; Rae et al., 2021). These limitations highlight the ongoing need for more adaptive and scale-aware methods to determine effective domain weights across various model sizes and target scenarios.
This work explicitly investigates and confirms the scale-dependence of optimal domain mixing, characterizing its scaling behavior. Based on these findings, we develop a practical methodology: optimizing domain weights at smaller, affordable scales and leveraging derived scaling laws to predict optimal mixing at much larger target scales. We lay out an overview of this work and main results in Fig. 1. Our contributions are summarized as follows.
1
Principled algorithmic framework for optimal domain mixing. Investigating the scaling law of domain mixing extends beyond mere empirical study. It requires a mathematical definition of optimal domain mixing and a tractable algorithm to solve for optimal weights. Our first contribution is formulating the optimal mixing problem as a bi-level optimization. However, existing general bi-level optimization techniques (Colson et al., 2007; Zhang et al., 2024) are intractable in this context due to their reliance on second-order information. We propose a novel approach tailored to our problem context that leverages scaling laws to estimate the dependence of the learned model’s loss on the weights, effectively reducing the bi-level problem to a single level. Our algorithm requires retraining models only linearly in the number of data domains, making it feasible for exploratory studies.
2
Uncovering and quantifying the scale-dependence of optimal domain composition.
Leveraging the algorithm developed in 1, we conduct empirical studies to optimize domain weights at different training data scales. Our results demonstrate that the optimal data composition varies with the scale of the training data, suggesting that the common practice of empirically determining an optimal composition using small-scale experiments will not yield optimal data mixtures for larger scales. We further derive an analytical framework for modeling the functional relationship between optimal data composition and training data scales.
3
Practical algorithm for optimal domain mixing. While the algorithm in 1 has made optimal domain mixing feasible for exploratory studies, its retraining requirements limit its practicality to smaller scales. To enable data composition optimization at large scales, we propose AutoScale. This method works by finding optimal data compositions at smaller, computationally feasible scales, fitting a predictor using our analytical model for the scale-dependency of optimal composition mentioned in 2, and finally using this predictor to determine optimal data composition at larger scales. Since one only needs to train models on small data scales where re-training is affordable, AutoScale does not require using proxy models with a smaller parameter size, avoiding transferability issues between domain weights optimized with different model sizes.
4
Robust performance gains across models and datasets. Our evaluation of AutoScale on both decoder-only and encoder-only models demonstrates its consistent ability to achieve significant computational savings. For instance, in pre-training GPT-2 Large (Radford et al., 2019) on the RedPajama dataset, AutoScale decreases validation perplexity 28% faster than any baseline, with up to 38% speed-up compared to training without reweighting. It also achieves the best overall performance across downstream tasks. Additionally, we present intriguing findings regarding the varying benefits of traditionally perceived high-quality and low-quality data sources across different training scales. Specifically, we observe that data sources with standardized formats, such as Wikipedia and scientific papers—often regarded as high-quality—are most beneficial at smaller scales but exhibit sharp diminishing returns as the training data scales up. Conversely, with increased compute, data sources containing diverse examples, such as CommonCrawl, demonstrate continued reductions in training loss even at considerably large training data scales.
Figure 1: Overview and main results. I. Optimizing domain weights with the proposed Direct Data Optimization (DDO) algorithm for pre-training 774M Decoder-only LMs (GPT-2 Large). Optimal domain weights depend on the scale of training data. A consistent shift can be observed. Using domain weights optimized for a different scale yields sub-optimal results, failing to fully realize the benefits of domain reweighting. II. Optimal domain data quantity (y-axis) for different training data scales (x-axis) shows high linearity ($R^{2}=0.998$) on log-log plot, suggesting the shifting pattern can be well predicted by exponential-style functions. We fit AutoScale to predict optimal domain weights for larger training scales. As we scale up, data sources with diverse samples (e.g., C4) are upweighted relative to domains with standard format (e.g., Wikipedia). III. Training 774M Decoder-only LMs for 10B tokens (96k steps). AutoScale-predicted domain weights decrease validation PPL at least $28\%$ faster than any baseline with up to $38\%$ speed up, achieving best overall task performance.

### v5

# AutoScale: Scale-Aware Data Mixing for Pre-Training LLMs

@@ Abstract
Domain reweighting is an emerging research area aimed at adjusting the relative weights of different data sources to improve the effectiveness and efficiency of LLM pre-training. We show that data mixtures that perform well at smaller scales may not retain their advantage at larger scales, challenging the existing practice of determining competitive mixtures in small-scale experiments and directly applying them at much larger scales. To address this, we propose AutoScale, a two-stage, scale-aware data composition framework. First, AutoScale fits a parametric model that predicts the model’s loss under different data compositions, then uses it to find an approximate best allocation at smaller, more manageable budgets. Next, leveraging a novel theoretical analysis of how optimal compositions evolve with scale, AutoScale extrapolates that composition to larger budgets without further retraining. Empirically, AutoScale accelerates convergence and improves downstream performance. For instance, when pre-training GPT-2 Large, it achieves a 28% faster perplexity reduction than baselines and up to a 38% speed-up over unweighted training, while yielding best-average results on various downstream tasks. Overall, our findings illustrate how domain importance shifts with training scale, underscoring the need for scale-dependent data curation in LLM training. Our code is open-sourced [^1] https://github.com/feiyang-k/AutoScale.
@@ 1 Introduction
Large language models (LLMs) are pre-trained on vast datasets sourced from diverse domains. However, the immense computational demands of this process, coupled with limited resources, create a pressing need to enhance the effectiveness and efficiency of pre-training. A promising approach to address this challenge is through domain reweighting—adjusting the ratio (or weights) of data from different sources.
However, developing a principled and efficient framework for determining an optimal data mix remains challenging. Many industry pipelines still rely on trial-and-error heuristics (Rae et al., 2021; Grattafiori et al., 2024) or reuse domain weights designed for previous models (Mehta et al., 2024), without a systematic approach for deciding how much of each domain to include. The seminal domain-optimization work by Xie et al. (2024) attempted to upweight “difficult” domains, but later work (Fan et al., 2023) reported instability and only limited validation-loss improvements, partly because the chosen optimization objective does not robustly align with the model’s ultimate test-time performance. Therefore, recent methods (Liu et al., 2024; Ye et al., 2024) focus on directly optimizing domain weights for lower validation loss. However, the highly complex relationship between domain weights and model performance makes such optimization expensive. A common strategy of these works is to reduce costs is to train multiple times at smaller scales, identify a “best” mix, and then assume it transfers to large-scale pre-training . Yet our experiments show that compositions found at smaller scales may not remain competitive when training is scaled up, whereas directly optimizing at full scale is infeasible. This yields a dilemma: either accept small-scale solutions that may not transfer or attempt large-scale optimization that is prohibitively costly.
To resolve this dilemma, we develop a novel theoretical analysis that shows how the optimal domain composition evolves at different scales. Building on this insight, we propose a two-stage framework, AutoScale, for scale-aware domain reweighting. In the first stage, we fit a parametric model that predicts the model’s loss under different data compositions, then use it to discover an approximate optimal allocation at smaller, more manageable budgets. Next, we apply our theoretical result to extrapolate that allocation to larger budgets—without re-optimizing at full scale—thus bridging the gap between small-scale optimization and full-scale pre-training.
Our experiments, conducted on both decoder-only and encoder-only architectures, consistently that AutoScale speeds up convergence and yields favorable downstream task performance. For instance, in pre-training GPT-2 Large on the RedPajama dataset, our approach achieves a 28% faster perplexity reduction compared with any baseline and up to a 38% speed-up over unweighted training, while also delivering the best downstream-task performance. Moreover, we made surprising empirical observations that data sources traditionally viewed as “high-quality” (e.g., Wikipedia and scientific papers) excel at smaller scales but exhibit sharp diminishing returns as the training grows. Meanwhile, domains containing more diverse examples (e.g., CommonCrawl) continue offering loss reductions at larger scales, underscoring the importance of scale-aware data curation.